40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Sat Nov 12 01:09:56 2016, gautam, Update, LSC, Recovering DRMI locking 
    Reply  Wed Nov 16 03:10:01 2016, gautam, Update, LSC, DRMI locked on 1f and 3f signals 
       Reply  Wed Nov 16 08:14:43 2016, Steve, Update, LSC, DRMI locked on 1f and 3f signals 5hrs.png
       Reply  Mon Nov 21 14:02:32 2016, gautam, Update, LSC, DRMI locked on 3f signals, arms held on ALS DRMIArms_Nov20.pdf
    Reply  Wed Nov 23 16:21:02 2016, gautam, Update, LSC, ITMY UL glitches are back ITMYproblemsBack.png
       Reply  Mon Nov 28 10:27:13 2016, gautam, Update, SUS, ITMY UL glitches are back ITMY_satboxSpectra.pdfITMY_UL_testerBox.png
       Reply  Tue Nov 29 11:07:37 2016, Steve, Update, LSC, ITMY UL glitches are back gliching400d.png
          Reply  Wed Nov 30 01:47:56 2016, gautam, Update, LSC, Suspension woes IMCwoes.png
             Reply  Wed Nov 30 17:08:56 2016, gautam, Update, LSC, Binary output breakout box removed 
                Reply  Thu Dec 1 02:19:13 2016, gautam, Update, LSC, Binary output breakout box restored MC1_glitch_fast.pngITMY_testerGlitch.pngITMYsatbox.pngULcomparison.pdf
                   Reply  Thu Dec 1 08:02:57 2016, Steve, Update, LSC, glitching ITMY_UL_LL ITMY_UL_LL.png
                      Reply  Fri Dec 2 11:56:42 2016, gautam, Update, LSC, MC1 LEMO jiggled IMG_3474.JPGMC1_LR_dead.png
                         Reply  Mon Dec 5 15:05:37 2016, gautam, Update, LSC, MC1 glitches are back Mc1glitches.png
                      Reply  Thu Dec 8 10:13:43 2016, Steve, Update, LSC, glitching ITMY_UL has a history glitching__ITMY-UL_2007.png
Message ID: 12619     Entry time: Wed Nov 16 03:10:01 2016     In reply to: 12611     Reply to this: 12620   12630
Author: gautam 
Type: Update 
Category: LSC 
Subject: DRMI locked on 1f and 3f signals 

After much trial and error with whitening gains, demod phases and overall loop gains, I was finally able to lock the DRMI on both 1f and 3f signals! I went through things in the following order tonight:

  1. Lock the arms, dither align
  2. Lock the PRMI on carrier and dither align the PRM to get good alignment
  3. Tried to lock the DRMI on 1f signals - this took a while. I realized the reason I had little to no success with this over the last few days was because I did not turn off the automatic unwhitening filter triggering on the demod screens. I had to tweak the SRM alignment while looking at the AS camera, and also adjust the demod phases for AS55 (MICH is on AS55Q) and REFL55 (SRCL is on REFL55I). Once I was able to get locks of a few seconds, I used the UGF servos to set the overall loop gain for MICH, PRCL and SRCL, after which I was able to revert the filter triggering to the usual settings
  4. Once I adjusted the overall gains and demod phases, the DRMI locks were very stable - I left a lock alone for ~20mins, and then took loop shape measurements for all 3 loops
  5. Then I decided to try transfering to 3f signals - I first averaged the IN1s to the 'B' channels for the 3 vertex DOFs using cds avg while locked on the 1f signals. I then set a ramp time of 5 seconds and turned the gain of the 'A' channels to 0 and 'B' channels to 1. The transition wasn't smooth in that the lock was broken but was reacquired in a couple of seconds.
  6. The lock on 3f signals was also pretty stable, the current one has been going for >10 minutes and even when it loses lock, it is able to reacquire in a few seconds

I have noted all the settings I used tonight, I will post them tomorrow. I was planning to try a DRFPMI lock if I was successful with the DRMI earlier tonight, but I'm calling it a night for now. But I think the DRMI locking is now back to a reliable level, and we can push ahead with the full IFO lock...

It remains to update the auto-configure scripts to restore the optimized settings from tonight, I am leaving this to tomorrow as well...


Updated 16 Nov 2016 1130am

Settings used were as follows:

1f/3f DOF Error signal Whitening gain (dB) Demod phase (deg) Loop gain Trigger
DRMI Locking 16 Nov 2016
1f MICH (A) AS55Q 0 -42 -0.026 POP22I=1
1f PRCL (A) REFL11I 18 18 -0.0029 POP22I=1
1f SRCL (A) REFL55I 18 -175 -0.035 POP22I=10
3f MICH (B) REFL165Q 24 -86 -0.026 POP22I=1
3f PRCL (B) REFL33I 30 136 -0.0029 POP22I=1
3f SRCL (B) REFL165I and REFL33I - - -0.035 POP22I=10

 

ELOG V3.1.3-